# Real-time Photorealistic Dynamic Scene Representation and Rendering with 4D Gaussian Splatting

### 18 Jan 2024 v2

@@ 2 Related work
Our method is also inspired by Kerbl et al. (2023). Critically, going beyond 3D Gaussians, we formulate a scene primitive, termed 4D Gaussians, to represent the dynamic scenes. This is more than the baseline extension that assumes dynamic scenes are generated by a fixed set of 3D Gaussians. This design allows us to end-to-end optimize the model, instead of otherwise the inefficient and less effective frame-by-frame training, in addition to the merit of eliminating the need for any prior regularization. Consequently, our approach can flexibly deal with the new object entering the scene, which is difficult to handle by the dynamic 3D Gaussians. Moreover, while dynamic 3D Gaussians omit view-dependent colors to avoid interference for tracking relationships, this compromises their ability to model non-Lambertian effects. In contrast, our method not only allows the appearance of Gaussian representations to vary with the view direction but also permits their view-dependent colors to evolve over time. Thus, our approach is more general and versatile to handle complex scenes and challenging monocular videos in real-world applications.
@@ 4.2 Implementation details
To assess the versatility of our approach, we did not extensively fine-tune the training schedule across different datasets. By default, we conducted training with a total of 30,000 iterations, a batch size of 8, and halted densification at the midpoint of the schedule. We adopted the settings of Kerbl et al. (2023) for hyperparameters such as loss weight, learning rate, and threshold. At the outset of training, we initialized both $q_{l}$ and $q_{r}$ as unit quaternions to establish identity rotations and set the initial time scaling to half of the scene’s duration. While the 4D Gaussian theoretically extends infinitely, we applied a Gaussian filter with marginal $p(t)<0.05$ when rendering the view at time $t$. For scenes in the Plenoptic Video dataset, we further initialized the Gaussian with 100,000 extra points distributed uniformly on the sphere encompassing the entire scene to fit the distant background that colmap failed to reconstruct and terminate its optimization after 10,000 iterations. Following the previous work, the LPIPS (Zhang et al., 2018) in the Plenoptic Video dataset and the D-NeRF dataset are computed using AlexNet (Krizhevsky et al., 2012) and VGGNet (Simonyan & Zisserman, 2014) respectively.

### 22 Feb 2024 v3

@@ 2 Related work
Recently, there have been substantial efforts (Chen & Wang, 2024) in extending 3D Gaussian Splatting into dynamic scenes. Beyond the pioneer work Luiten et al. (2024) mentioned above, Yang et al. (2023); Wu et al. (2023); Liang et al. (2023) propose to model the geometry and dynamic of scenes by the joint optimization of Gaussians in canonical space and a deformation field. (Kratimenos et al., 2023) encourages locality and rigidity between points by factorizing the motion in the scene into a few neural trajectories. These works ingeniously incorporate a prior of topological invariance into their representations, making them well-suited for reconstructing dynamic scenes from monocular videos. However, they assume that dynamic scenes are generated by a fixed set of 3D Gaussians and the elements composing the scene are always visible. In contrast, by formulating a novel 4D scene primitive, we discard their underlying assumptions and circumvent the need to maintain ambiguous and complex tracking relationships, thereby facilitating a more flexible and versatile approach to handling complex scenes in real-world applications.
@@ 4.2 Implementation details
To assess the versatility of our approach, we did not extensively fine-tune the training schedule across different datasets. By default, we conducted training with a total of 30,000 iterations, a batch size of 4, and halted densification at the midpoint of the schedule. We adopted the settings of Kerbl et al. (2023) for hyperparameters such as loss weight, learning rate, and threshold. At the outset of training, we initialized both $q_{l}$ and $q_{r}$ as unit quaternions to establish identity rotations and set the initial time scaling to half of the scene’s duration. While the 4D Gaussian theoretically extends infinitely, we applied a Gaussian filter with marginal $p(t)<0.05$ when rendering the view at time $t$. For scenes in the Plenoptic Video dataset, we further initialized the Gaussian with 100,000 extra points distributed uniformly on the sphere encompassing the entire scene to fit the distant background that colmap failed to reconstruct and terminate its optimization after 10,000 iterations. Following the previous work, the LPIPS (Zhang et al., 2018) in the Plenoptic Video dataset and the D-NeRF dataset are computed using AlexNet (Krizhevsky et al., 2012) and VGGNet (Simonyan & Zisserman, 2014) respectively.